The equity section of Cyril Corporation’s balance sheet shows the following. Preferred stock—5% cumulative, $25 par value, 10,000 shares issued and outstanding $ 250,000 Common stock—$10 par value, 45,000 shares issued and outstanding 450,000 Retained earnings 267,500 Total stockholders’ equity $ 967,500

Determine the book value per share of common stock under two separate situations.

1. No preferred dividends are in arrears at the current date.
2. Three years of preferred dividends are in arrears at the current date.

Determine the book value per share of the preferred and common stock for no preferred dividends are in arrears.

Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to common shares / Number of common shares outstanding = Book value per common share
x / 45,000 =

Determine the book value per share of the preferred and common stock for three years of preferred dividends are in arrears.

Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to commont / Number of common shares outstanding = Book value per common share
x / 45,000 =

I am having a hard time finding the values for x in both charts.
Answer #1
1. No preferred dividends are in arrears at the current date.
Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to common shares / Number of common shares outstanding = Book Value per Common Share
717500 / 45000 = 15.94
2. Three years of preferred dividends are in arrears at the current date.
Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to common shares / Number of common shares outstanding = Book Value per Common Share
680000 / 45000 = 15.11

Figures shown above have been calculated in the following manner:

1. No preferred dividends are in arrears at the current date.
Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to common shares / Number of common shares outstanding = Book Value per Common Share
=967500-250000 / 45000 = 717500/45000
2. Three years of preferred dividends are in arrears at the current date.
Book Value per Common Share
Choose Numerator: / Choose Denominator: = Book Value per Common Share
Stockholders' equity applicable to common shares / Number of common shares outstanding = Book Value per Common Share
=967500-250000-(250000*5%*3) / 45000 = 680000/45000